Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

2021-02-09

Observers:  Maxime Deckers, Simon Prentice
Support: Jacco Terwel, Mark Magee
TAT: Cristina Barbarino, Giuliano Pignata, Tomás E. Müller Bravo

EFOSC Calibrations: bias, calib


Weather:
Start: humidity 36, wind 6.8, seeing 0.8
Middle: humidity 40, wind 5.3, seeing 0.9
End: humidity 34, wind 2.0, seeing 0.6

Comments:
Some time lost at the start of the night as there was no day-time operator present to perform calibrations. 
Moreover, due to technical problems the re-acquisitions took longer than expected.
Final standard missed due to duplicate gr11 spectrum.

Started with EFOSC

 UT Target Type Set up
 Slit Exposure Airmass Seeing Notes
 01:05
 AT2021blz  Follow up
 gr13 1.0
 1800s
 1.0
 1.2
 01:54 GD71 Standard gr11, 13, 16 1.0+1.5 100s 1.6 0.84
 02:15 SN2020znr Follow Up gr11, gr16 1.0 1800s 2.5 1.2 
 03:43 ZTF20acvjagm Follow up gr13 1.0 1800s 1.4 0.65 
 04:22 AT2020xys Follow up gr11, gr16 1.0 1500s 1.2 0.77 
 05:28 LTT3864 Standard gr11, 13, 16 1.0+1.5 100s 1.0 0.85 
 05:50 AT2021bnw     Follow up gr11, gr16 1.0 1800 x3, 1800 x2 1.4 0.85 Communication error led to 3 x gr11
         
Comments